Montserrat Volcano Observatory

Daily Report
Report for the period 6 pm 14 February to 6 pm 15 February 1998

Activity at the volcano was about the same as yesterday with dust levels low and rockfall events dominant.

Between 16:00 yesterday and 16:00 today there were 5 hybrid earthquakes, 13 long period earthquakes, 1 volcano-tectonic earthquake and 100 distinct rockfall signals. One of the long period earthquakes preceded a rockfall. There were three periods of elevated seismic activity during which most of the rockfalls occurred. These were experienced between 5:30 pm and 7 pm last night, between midnight and 2 am this morning and between 10 am to 12 noon today.

The periods of heightened activity correlated with vigorous ash venting and steam plumes. Views of the dome from the police boat late this morning showed that there were several large rockfalls down the White River, but none of them reached further than the base of the talus slope below the Boxing Day dome. Dark ash clouds drifted westwards from the volcano. MVO would like to thank the crew of the police boat for their assistance today.

There was a further collaborative GPS survey today between MVO and the University of Puerto Rico. Stations were occupied at Harris, Lookout, St. George's Hill, Perche's Mountain and Lee's Yard. The results of these measurements are currently being processed.

Dust levels at all of the monitoring sites remained low today.

Dr. Sue Loughlin of the British Geological Survey arrived yesterday to take over duties as deputy chief scientist. Christian Antenor-Habazac from the Volcano Observatory in Guadeloupe left today after a two week visit to help with installation of new electronic equipment.

Montserrat Volcano Observatory